Copper has a specific heat of 0.382 J/g×°C. The temperature of an unknown mass of copper increases by 4.50°C when it absorbs 3.97 J of heat. What is the mass of the copper?

1
Expert's answer
2021-03-24T01:09:33-0400

Given,

Specific heat of copper "c=0.382J/gc^{\circ}"

Increament in temprature "\Delta T=4.5^{\circ}"

Heat absorbs Q=3.97 J


As we know,

"Q=mc\Delta T"


"\Rightarrow 3.97=m\times 0.382\times 4.5\\\Rightarrow m=\dfrac{3.97}{1.719}=2.309g"


Hence The mass of the copper is 2.309g